# Simplot: Enhancing Chart Question Answering by Distilling Essentials

### v2

@@ Abstract
Recently, interpreting complex charts with logical reasoning has emerged as challenges due to the development of vision-language models. A prior state-of-the-art (SOTA) model has presented an end-to-end method that leverages the vision-language model to convert charts into table format utilizing Large Language Model (LLM) for reasoning. However, unlike natural images, charts contain a mix of essential and irrelevant information required for chart reasoning, and we discover that this characteristic can lower the performance of chart-to-table extraction. In this paper, we introduce Simplot, a method designed to extract only the elements necessary for chart reasoning. The proposed method involves two steps: 1) training to mimic a simple plot that contains only the essential information from a complex chart for table extraction, followed by 2) performing reasoning based on the table. Our model enables accurate chart reasoning without the need for additional annotations or datasets, and its effectiveness is demonstrated through various experiments. Furthermore, we propose a novel prompt mimicking how human interpret charts for more accurate reasoning. Our source code is available at https://github.com/sangwu99/Simplot.
@@
Wonjoong Kim1thanks: These authors contributed equally.,  Sangwu Park2111The authors also provide pre-trained model for chart to table extraction. We use this model to extract the table and conduct reasoning in the same setting of Deplot and Simplot.,  Yeonjun In1,  Seokwon Han1,  Chanyoung Park1thanks: Corresponding author. 1KAIST   2Chung-Ang University {wjkim, yeonjun.in, dulwich10, cy.park}@kaist.ac.kr  sangwu99@cau.ac.kr
@@ 2.1 Preprocessing Stage
1) Simple Chart Generation. We generate a simple chart by excluding irrelevant information from the original chart, keeping only essential elements required for reasoning. Specifically, since each of the original charts in the dataset is annotated with a table in the CSV format, we use a Python library (i.e., Matplotlib) to plot a chart based on the table. The chart generated in this manner is considered as a simplified version of the original chart, i.e., simple chart. This process requires no separate training or additional costs, and it can be executed offline by running a simple code snippet.
2) Row-Column Rendering. Inspired by rendering questions over images used for QA (Lee et al., 2023), we aim to improve the table extraction accuracy by rendering information about the rows and columns that should be included in the table onto the image, enabling the model to utilize this information to extract more accurate tables when converting charts to tables. Note that in the training stage, since we are given the ground-truth chart-table pairs, we simply render rows and columns of each table onto its paired image containing the chart. However, it would not be feasible in the inference stage, since we would be only given the charts without tables. Instead, in the inference stage, we utilize a LMM to extract rows and columns from the chart and render them onto the image. Despite the relative low performance of LMM such as GPT-4 (Achiam et al., 2023) in reasoning about charts, we found they can accurately extract row and column information from charts, since such information is given as text and structured in a relatively simple manner. Detailed description of the process is presented in Appendix B.
@@ 2.2 Training Stage: Chart-to-Table Extraction
Our proposed chart-to-table extraction approach consists of two phases: Phase 1) Training a teacher encoder and a table decoder by performing the chart-to-table extraction given simple charts containing only the essential information for reasoning, rather than complex original charts; Phase 2) Training a student encoder via contrastive learning by extracting the table given the original chart, while being distilled the knowledge to embed original charts to the embedding space of simple charts.
@@ 2.2.2 Phase 2: Encoding Original Chart into Representation Space of Simple Chart
As shown in Fig. 2, after training the teacher encoder $E_{chart}^{teacher}$ with simple charts as inputs, our goal is to train a student encoder $E_{chart}^{student}$ in a way that the representation of an original chart image obtained from $E_{chart}^{student}$ closely matches the representation of its corresponding simplified chart obtained from $E_{chart}^{teacher}$. This is to enable $E_{chart}^{student}$ to encode any original chart into the representation space of simplified chart images, thereby improving the accuracy of table generation. Then, given the chart representations, we generate table representations ${z}$ using the frozen table decoder $D_{table}$ that is trained in Phase 1. More precisely, given an original chart $A$ and the simplified chart $P$ as a positive sample, we generate a negative sample $N$ by randomly shuffling values in the simple chart $P$. Then, we define a triplet loss using the representations of the original chart $z_{a}$, its corresponding positive sample $z_{p}$, and negative sample $z_{n}$ obtained from the sequentially passing $A,P$ and $N$ into chart encoders and the table decoder $D_{table}$ as follows:
@@
where $d$ denotes the distance defined as $d(z_{i},z_{j})=\|z_{i}-z_{j}\|_{2}$, and $m$ denotes the margin which induces the anchor $z_{a}$ to be closer to the positive sample $z_{p}$ and farther from the negative sample $z_{n}$. Through this process, the student chart encoder $E_{chart}^{student}$ is trained to extract only the essential information from original charts, even if they contain information that is irrelevant to chart reasoning. By optimizing Eq. 2.2.2, $E_{chart}^{student}$ focuses on subtle differences in the representations of positive and negative samples that would be highly similar in the representation space, which is expected to enhance the precise value mapping for table generation.
The representation $z_{a}$ obtained after passing the original chart $A$ through the student encoder $E_{chart}^{student}$ and the table decoder $D_{table}$ is then fed into the frozen fully connected layer ($FC$) trained in Phase 1 to generate a table.
@@ 2.3 Inference Stage: Reasoning with Extracted Table
Human-oriented chart instruction. Furthermore, we present a novel prompt that is specifically designed for chart reasoning, named Human-oriented chart instruction. Although prompt engineering is a widely studied topic for processing natural images (Kim et al., 2023; Li et al., 2022), it is crucial to develop specialized prompts specifically tailored for charts, due to the inherent nature of chart reasoning tasks, i.e., they require more advanced advanced thinking steps for interpretation. Our proposed prompt provides instructions that mimic how humans interpret charts, enabling more accurate reasoning within the LMM. To our best knowledge, this is the first work to study a prompt specifically targeting charts, and its effectiveness will be demonstrated in Section 3.1.2. We provide a snippet of our proposed prompt in Fig. 4. For detailed description of Human-oriented chart instruction, please refer to Appendix C.
@@ 3 Experiments
Baselines. In this paper, we use various models as baselines that can perform chart reasoning. The baselines we employ are divided into three categories including Vision-Language pre-trained models (VLP), fully supervised models for chart reasoning, and models which extract table and perform reasoning using LLMs. Detailed descriptions of the models are provided in the Appendix D.2.
@@
Furthermore, we propose a proper metric to evaluate the chart extraction performance, employing a slightly modified metric derived from Relative Mapping Similarity ($RMS$), presented by Deplot (Liu et al., 2022a). We discover that the vanilla $RMS$ does not accurately measure the chart-to-table performance in certain cases, and thus we introduce a metric named Relative Distance ($RD$). For the differences between $RMS$ and $RD$, please refer to the Appendix D.3.
@@ 3.1.2 Results on Chart Reasoning
The performance of question answering on the ChartQA dataset is summarized in Table 2. 1) In general, we observe that methods designed with a focus on chart reasoning (i.e., ‘Supervised’ and ‘Table’) outperform VLP in QA tasks. This demonstrates that despite the superior performance of vision-language models in performing a wide range of tasks in the traditional natural image domain, they face difficulties in interpreting charts due to the significantly different characteristics of natural images and charts. 2) Furthermore, noticeable performance difference is observed among models aimed at chart reasoning. Methods using tables such as Deplot, Unichart, and Simplot, generally outperform models that conduct reasoning based solely on images (i.e., ‘Supervised’). This suggests that going through tables allows for more detailed reasoning, aligning well with our motivation that emphasizes the importance of effectively extracting tables. Additionally, models utilizing extracted tables and LMM can efficiently handle various tasks such as text summarization without any additional data and training. 3) Among the table-based reasoning models, Simplot performs the best by utilizing textual information, which is essential for obtaining the most accurate table, and by employing prompts mimicking the human reasoning process. Additional experimental results on PlotQA dataset and effectiveness of our method are presented in Appendix E and 3.2, respectively.
@@ 3.2 Further Analysis
Table 3: Ablation studies on each component of Simplot for table extraction and QA tasks.
@@
Deplot + Simplot 76.70 88.42 82.56
Performance on more Challenging Questions. It is worth noting that ‘Deplot+img+prompt’ now performs competitively to Simplot. However, we argue that this is mainly due to the simplicity of the ChartQA dataset, which makes the dataset insufficient to evaluate the table extraction capability of Simplot. For example, ChartQA contains questions that can be answered by referencing just one row of the extracted table.
To evaluate the models with more challenging questions, we randomly sample 100 test images and then provide GPT-4 with the following instruction: "Create a challenging question-answer pair that requires referencing at least two rows and two columns to solve.". Through this process, we generated 100 question-answer pairs and eventually obtained 85 pairs after manually filtering out inaccurate pairs, and then evaluate ‘Deplot+img+prompt’ and Simplot on these question-answer pairs (lower part of Table 5). We observe that the performance gap greatly increases as the questions get more challenging, which demonstrates the importance of extracting accurate tables when encountering challenging questions. For the specific example of the challenging question and how the extracted tables from Simplot and Deplot lead to differences in complex QA performance, please refer to Appendix I.
Table 5: Effectiveness of proposed prompt on ChartQA.
@@
Case Studies. We provide case studies showing how Simplot address the concerns we have raised, compared to our main baseline, Deplot (Liu et al., 2022a).
Firstly, we verify whether Simplot correctly captures textual information within charts. Fig. 7 illustrates that the table extracted by Deplot fails to consider the counting unit, i.e., "million/billion," present in the chart, while Simplot correctly captures the textual information.
We also confirm that rendering columns and rows of a table onto the image is beneficial for extracting an accurate table. In Fig. 8, Deplot (Liu et al., 2022a) omits a relevant row (i.e., "Montenegro"), redundantly extracts several rows, and even generates entirely irrelevant rows such as "CC BY," which would result in a poor performance in reasoning tasks. In contrast, Simplot accurately generates a table by extracting only the relevant rows, ensuring accurate reasoning.
@@
Furthermore, we demonstrate the effectiveness of Human-oriented chart instruction. Fig. 6 illustrates the difference between interpreting the same chart with/without using the proposed prompt. While the model fails to produce an accurate answer without the prompt, the model generates a correct answer in a process similar to how humans interpret charts when using the prompt. This highlights the effectiveness of our prompt designed to mimic human reasoning. Please refer to Appendix H for further case studies and Appendix J for error analysis.
@@ 5 Conclusion
We propose Simplot, which extracts only the essential information required for chart reasoning. Furthermore, by leveraging textual information inside the charts, Simplot enables accurate chart reasoning. Finally, we propose a novel prompt specifically designed for chart reasoning. It provides instruction to LMM imitating how humans interpret chart, enabling more precise reasoning. Through extensive experiments, we demonstrate that Simplot effectively handle concerns raised by existing works while improving performance, and also can be applied to other existing table-based reasoning model.
@@
Despite the outstanding performance of Simplot in chart-to-table extraction, there is room for improvement. During the process of rendering columns and rows onto the images, inaccuracies in extracting columns and rows can lead to failures in effectively converting charts into tables. Particularly, it can be challenging to extract appropriate columns and rows when the texts in the chart are complex. We provide a detailed error analysis related to this issue in the Appendix J. To address these problems, future work could involve verifying and ensuring that the extracted rows and columns maintain a coherent context, aiming for more accurate rendering and extraction.
@@ Appendix A Complete Related Work
Radford et al. (2021) and Jia et al. (2021) facilitated zero-shot adaptation utilizing contrastive learning for various tasks on multimodal data. However, their drawback resides in their limitation to generate language, diminishing their suitability for open-ended tasks such as VQA.
@@ D.2 Compared Methods
The third category consists of table-based reasoning models, including Deplot (Liu et al., 2022a) and Unichart (Masry et al., 2023). Deplot converts charts to tables similar to Simplot and employs prompt-based methodologies such as Chain-of-Thought (CoT) (Wei et al., 2022), Program-of-Thought (PoT) (Chen et al., 2022a) alongside using an LLM for question reasoning. Although Unichart is trained to perform QA directly from charts, it is also pre-trained to generate tables. In this paper, we use this variant of Unichart to extract tables and apply it in the same setting as Deplot.

### v3

@@ Abstract
Recently, interpreting complex charts with logical reasoning has emerged as challenges due to the development of vision-language models. A prior state-of-the-art (SOTA) model has presented an end-to-end method that leverages the vision-language model to convert charts into table format utilizing Large Language Model (LLM) for reasoning. However, unlike natural images, charts contain a mix of essential and irrelevant information required for chart reasoning, and we discover that this characteristic can lower the performance of chart-to-table extraction. In this paper, we introduce Simplot, a method designed to extract only the elements necessary for chart reasoning. The proposed method involves two steps: 1) training to mimic a simple plot that contains only the essential information from a complex chart for table extraction, followed by 2) performing reasoning based on the table. Our model enables accurate chart reasoning without the need for additional annotations or datasets, and its effectiveness is demonstrated through various experiments [^1] Our source code is available at https://github.com/sangwu99/Simplot..
@@
Wonjoong Kimthanks: These authors contributed equally.,  Sangwu Park11footnotemark: 1,  Yeonjun In,  Seokwon Han,  Chanyoung Parkthanks: Corresponding author. KAIST {wjkim, sangwu.park, yeonjun.in, dulwich10, cy.park}@kaist.ac.kr
@@ 2.1 Preprocessing Stage
1) Simple Chart Generation. We generate a simple chart by excluding irrelevant information from the original chart, keeping only essential elements required for reasoning. Specifically, since each of the original charts in the dataset is annotated with a table in the CSV format, we use a Python library (i.e., Matplotlib) to plot a chart based on the table. The chart generated in this manner is considered as a simplified version of the original chart, i.e., simple chart. This process requires no separate training or additional costs and it can be executed offline by running a simple code snippet, so it is scalable and generalizable across various datasets.
2) Row-Column Rendering. Inspired by rendering questions over images used for QA (Lee et al., 2023), we aim to improve the table extraction accuracy by rendering information about the rows and columns that should be included in the table onto the image, enabling the model to utilize this information to extract more accurate tables when converting charts to tables. Note that in the training stage, since we are given the ground-truth chart-table pairs, we simply render rows and columns of each table onto its paired image containing the chart. However, it would not be feasible in the inference stage, since we would be only given the charts without tables. Instead, in the inference stage, we utilize a LMM to extract rows and columns from the chart and render them onto the image. Despite the relative low performance of LMM such as GPT-4 (Achiam et al., 2023) in chart reasoning, we found they can accurately extract row and column information from charts, since such information is given as text and structured in a relatively simple manner. Detailed description of the process is presented in Appendix B.
@@ 2.2 Training Stage: Chart-to-Table Extraction
Our proposed chart-to-table extraction approach consists of two phases: Phase 1) Training a teacher encoder and a table decoder by performing the chart-to-table extraction given simple charts containing only the essential information for reasoning, rather than complex original charts; Phase 2) Training a student encoder by extracting the table given the original chart, while being distilled the knowledge by teacher encoder to embed original charts to the embedding space of simple charts.
@@ 2.2.2 Phase 2: Encoding Original Chart into Representation Space of Simple Chart
As shown in Fig. 2, after training the teacher encoder $E_{chart}^{teacher}$ with simple charts as inputs, our goal is to train a student encoder $E_{chart}^{student}$ in a way that the representation of an original chart image obtained from $E_{chart}^{student}$ closely matches the representation of its corresponding simplified chart obtained from $E_{chart}^{teacher}$. This is to enable $E_{chart}^{student}$ to encode any original chart into the representation space of simplified chart images, thereby improving the accuracy of table generation. Then, given the chart representations, we generate table representations ${z}$ using the frozen table decoder $D_{table}$ trained in Phase 1.
More precisely, given an original chart $A$ and the simplified chart $P$, we obtain the representations of the original chart $z_{a}$ and its corresponding sample $z_{p}$ from sequentially passing into chart encoders and the table decoder. For efficient learning, we define the triplet loss by adding a negative sample $N$, which is created by randomly shuffling the values of the simple chart $P$ and its representation $z_{n}$ as follows:
@@
where $d$ denotes the distance defined as $d(z_{i},z_{j})=\|z_{i}-z_{j}\|_{2}$, and $m$ denotes the margin which induces the anchor $z_{a}$ to be closer to the positive sample $z_{p}$ and farther from the negative sample $z_{n}$. The purpose of utilizing negative samples is to more precisely discriminate between data that may be mapped closely in the representation space because the charts are similar, and further details and effectiveness are presented in Appendix H.
Through this process, the student chart encoder $E_{chart}^{student}$ is trained to extract only the essential information from original charts, even if they contain information that is irrelevant to chart reasoning. The representation $z_{a}$ obtained after passing the original chart $A$ through the student encoder $E_{chart}^{student}$ and the table decoder $D_{table}$ is then fed into the frozen fully connected layer ($FC$) trained in Phase 1 to generate a table.
@@ 2.3 Inference Stage: Reasoning with Extracted Table
Human-oriented chart instruction. Furthermore, we present a novel prompt that is specifically designed for chart reasoning, named Human-oriented chart instruction. Although prompt engineering is a widely studied topic for processing natural images (Kim et al., 2023; Li et al., 2022), it is crucial to develop specialized prompts specifically tailored for charts, due to the inherent nature of chart reasoning tasks, i.e., they require more advanced advanced thinking steps for interpretation. Our proposed prompt provides instructions that mimic how humans interpret charts, enabling more accurate reasoning within the LMM and its effectiveness will be demonstrated in Section 3.1.2. We provide a snippet of our proposed prompt in Fig. 4. For detailed description of Human-oriented chart instruction, please refer to Appendix C.
@@ 3 Experiments
Baselines. In this paper, we use various models as baselines that can perform chart reasoning. The baselines we employ are divided into three categories including Vision-Language pre-trained models (VLP), fully supervised models for chart reasoning, and models which extract table and perform reasoning using LLMs. For a fair comparison, we employ GPT-4V for all methods utilizing extracted table. Detailed descriptions of the models are provided in the Appendix D.2.
@@
Furthermore, we propose a proper metric to evaluate the chart-to-table extraction performance, employing a slightly modified metric derived from Relative Mapping Similarity ($RMS$), presented by Deplot (Liu et al., 2022a). We discover that the vanilla $RMS$ does not accurately measure the chart-to-table performance in certain cases, and thus we introduce a metric named Relative Distance ($RD$). For the differences between $RMS$ and $RD$, please refer to the Appendix D.3.
@@ 3.1.1 Results on Chart-to-Table Extraction
We also have mentioned that GPT-4V can achieve more accurate text extraction rather than numerical extraction, which is why we utilize it for row and column extraction. Nevertheless, to further validate the necessity of a model specifically targeting charts, we also conduct table extraction (both text and numerical information) using GPT-4V, and the results are presented in the Table 1. As previously argued, GPT-4V is well-suited for extracting text information from charts, so it demonstrated comparable performance on pie charts where values and rows/columns are primarily given as text. However, for line charts, where most values must be inferred from the image rather than being provided as text, GPT-4V showed significantly lower performance compared to other models.
This explains the low QA performance in Table 2 as well, and conducting QA using tables extracted by GPT-4V is inefficient both in terms of performance and cost. Furthermore, these results support our argument that existing vision-language models still lack effectiveness when dealing with chart images while performing well on natural images, so methods specifically trained to target charts are necessary.
@@ 3.1.2 Results on Chart Reasoning
The performance of question answering on the ChartQA dataset is summarized in Table 2. 1) In general, we observe that methods designed with a focus on chart reasoning (i.e., ‘Supervised’ and ‘Table’) outperform VLP in QA tasks. This demonstrates that despite the superior performance of vision-language models in performing a wide range of tasks in the traditional natural image domain, they face difficulties in interpreting charts due to the significantly different characteristics of natural images and charts. 2) Furthermore, noticeable performance difference is observed among models aimed at chart reasoning. Methods using tables such as Deplot, Unichart, and Simplot, generally outperform models that conduct reasoning based solely on images (i.e., ‘Supervised’). This suggests that going through tables allows for more detailed reasoning, aligning well with our motivation that emphasizes the importance of effectively extracting tables. Additionally, models utilizing extracted tables and LMM can efficiently handle various tasks such as OpenCQA (Kantharaj et al., 2022a) without any additional data and training. 3) Among the table-based reasoning models, Simplot performs the best by utilizing textual information, which is essential for obtaining the most accurate table, and by employing prompts mimicking the human reasoning process. Additional experimental results on PlotQA dataset and effectiveness of our method are presented in Appendix E and 3.2, respectively.
@@ 3.2 Further Analysis
Table 3: Ablation studies on each component of Simplot for chart-to-table extraction and QA tasks.
@@
Deplot + Simplot 78.07 88.42 83.24
Performance on more Challenging Settings. It is worth noting that ‘Deplot+img+prompt’ now performs competitively to Simplot. However, we argue that this is mainly due to the simplicity of the ChartQA dataset, which makes the dataset insufficient to evaluate the table extraction capability of Simplot. For example, ChartQA contains questions that can be answered by referencing just one row of the extracted table.
To evaluate the models with more challenging questions, we randomly sample 100 test images and then provide GPT-4 with the following instruction: "Create a challenging question-answer pair that requires referencing at least two rows and two columns to solve.". Through this process, we generated 100 question-answer pairs and eventually obtained 85 pairs after manually filtering out inaccurate pairs, and then evaluate ‘Deplot+img+prompt’ and Simplot on these question-answer pairs (lower part of Table 5). We observe that the performance gap greatly increases as the questions get more challenging, which demonstrates the importance of extracting accurate tables when encountering challenging questions. For the specific example of the challenging question and how the extracted tables from Simplot and Deplot lead to differences in complex QA performance, please refer to Appendix J.
@@
Table 5: Performance on more challenging questions.
Models Com. (Easy) Com. (Hard) Sequential
Deplot+img.+prompt 57.50 50.00 26.80
Simplot 59.20 58.10 36.50
Table 6: Performance on MultiChartQA dataset.
Furthermore, questions that necessitate referring to multiple charts (Zhu et al., 2024) are expected to be significantly impacted by the accuracy of table extraction when it comes to improving QA performance. The additional experimental results are presented in the Table 6. On the comparative dataset, which requires simultaneously comparing values from two or more charts, and the sequential dataset, which requires finding new information in one chart based on information from another chart, our proposed method consistently outperforms Deplot. Since these tasks require examining multiple charts simultaneously, the results strongly support our claim that accurate table extraction leads to better performance on complex tasks. Moreover, the comparative dataset is divided into easy and hard categories, and the greater performance differences observed on harder questions indicate that even small improvements in table extraction accuracy can significantly impact accurate QA.
Adaptability of Chart-to-table Extraction to another task. As we have argued before, extracting tables and combining them with LMM not only improves the accuracy of QA but also offers the advantage of immediately being applicable to other tasks such as OpenCQA (Kantharaj et al., 2022a) without additional training.
Table 7 presents the BLEU score (Papineni et al., 2002) of a model fine-tuned for OpenCQA (ChartInstruct), and those of models that utilize extracted tables (Deplot & Simplot). Here, we demonstrate that leveraging tables allows for effective performance on other tasks without additional training, while also proving that accurately extracting tables enhances accuracy when performing these tasks. Note that this dataset includes charts in a different format from those used to train Simplot, so training with the charts from this dataset could potentially improve performance.
Models BLEU score
Gemini 6.84
GPT-4V 3.31
ChartInstruct 5.27
Deplot 6.88
Simplot 7.11
Table 7: Open-ended question answering performance on OpenCQA dataset.
Case Studies. We verify whether Simplot correctly captures textual information within charts. Fig. 6 illustrates that the table extracted by Deplot fails to consider the counting unit, i.e., "million/billion," present in the chart, while Simplot correctly captures the textual information.
@@
Please refer to Appendix I for further case studies including the results that demonstrate the effectiveness of proposed prompt and Appendix K for error analysis.
@@ 5 Conclusion
We propose Simplot, which extracts only the essential information required for chart reasoning. Furthermore, by leveraging textual information inside the charts, Simplot enables accurate chart reasoning. We also propose a novel prompt specifically designed for chart reasoning, which guides LMM to imitate how humans interpret chart. Through extensive experiments, we demonstrate that Simplot effectively handle concerns raised by existing works while improving performance, and also can be applied to other various tasks.
@@
As with existing chart reasoning methods, Simplot is also expected to have limited performance on unseen out-of-distribution (OOD) data. This is a future research area that all chart reasoning methods, including ours, should consider. However, please note that our proposed method can address various downstream tasks with just one round of training to extract tables from OOD charts, unlike other methods that require separate training for each task. Meanwhile, we conduct inference without training for the OpenCQA dataset where no tables were provided, and we observed that Simplot outperformed the existing baselines, indicating that our proposed method ensures a certain level of generalization performance for OOD charts.
@@
Acknowledgements
This work was supported by Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2023-00216011, Development of artificial complex intelligence for conceptually understanding and inferring like human) and National Research Foundation of Korea(NRF) funded by Ministry of Science and ICT (NRF-2022M3J6A1063021).
@@ D.2 Compared Methods
  • ChartAssistant (Meng et al., 2024)
  • ChartInstruct (Masry et al., 2024)
LLM Textual info. Designed prompt
Unichart
Deplot
Simplot
Table 10: Comparison of models based on their utilization of LLM, textual information, and designed prompt.
The third category consists of table-based reasoning models, including Deplot (Liu et al., 2022a) and Unichart (Masry et al., 2023). Deplot converts charts to tables similar to Simplot and employs prompt-based methodologies such as Chain-of-Thought (CoT) (Wei et al., 2022), Program-of-Thought (PoT) (Chen et al., 2022a) alongside using an LLM for question reasoning. Although Unichart is trained to perform QA directly from charts, it is also pre-trained to generate tables. In this paper, we use this variant of Unichart to extract tables and utilize LLM to fairly compare with Deplot in the same setting.
Regardless of our application of Unichart in the same setting as Deplot, the original Unichart requires direct training for each downstream task. It has limitations in terms of adaptability and may exhibit suboptimal performance. Deplot and Simplot achieve performance gains and enhanced adaptability by performing inference through an LLM after chart-to-table extraction. Additionally, Simplot significantly improves the accuracy of table extraction and downstream tasks by leveraging textual information and utilizing prompts specifically designed for charts. The differences among the aforementioned models can be referred to Table 10.
As the existing chart reasoning models (Liu et al., 2022a; Masry et al., 2023, 2024; Meng et al., 2024; Liu et al., 2022b; Masry et al., 2022; Lee et al., 2023) require expensive computational time, they have reported the performance after a single run, which makes it difficult to conduct statistical comparison. Following the previous works, we only report the performance after a single run. However, we claim that our results improve the performance of chart question answering with significant margin.